Paynesville Corporation manufactures and sells a preservative used in food and drug manufacturing. The company carries no inventories. The master budget calls for the company to manufacture and sell 124,000 liters at a budgeted price of $255 per liter this year. The standard direct cost sheet for one liter of the preservative follows.

Direct materials (2 pounds @ $16) $ 32
Direct labor (0.5 hours @ $48) 24


Variable overhead is applied based on direct labor hours. The variable overhead rate is $140 per direct-labor hour. The fixed overhead rate (at the master budget level of activity) is $70 per unit. All non-manufacturing costs are fixed and are budgeted at $2.4 million for the coming year.

At the end of the year, the costs analyst reported that the sales activity variance for the year was $774,000 unfavorable.

The following is the actual income statement (in thousands of dollars) for the year.

Sales revenue $ 30,398
Less variable costs
Direct materials 3,298
Direct labor 2,810
Variable overhead 7,830
Total variable costs $ 13,938
Contribution margin $ 16,460
Less fixed costs
Fixed manufacturing overhead 1,170
Non-manufacturing costs 1,350
Total fixed costs $ 2,520
Operating profit $ 13,940


During the year, the company purchased 200,000 pounds of material and employed 52,400 hours of direct labor.


Required:

a. Compute the direct material price and efficiency variances.
b. Compute the direct labor price and efficiency variances.
c. Compute the variable overhead price and efficiency variances.

(For all requirements, enter your answers in whole dollars. Indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option.)
Answer #1

Solution a:

Budgeted contribution margin per unit = $255 - ($32 + $24 + 0.5*$140) = $129 per unit

Actual units sold = Budgeted units - (Unfavorable sales activity variance / Budgeted contribution margin per unit)

= 124000 - ($774,000 / $129) = 118000 units

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
200000 $16.49 $3,298,000.00 200000 $16.00 $3,200,000.00 236000 $16.00 $3,776,000.00
$98,000.00 U $576,000.00 F
Direct Material Price Variance Direct Material efficiency variance
Direct material price variance $98,000.00 U
Direct material efficiency variance $576,000.00 F

Solution b:

Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
52400 $53.63 $2,810,000.00 52400 $48.00 $2,515,200.00 59000 $48.00 $2,832,000.00
$294,800.00 U $316,800.00 F
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $294,800.00 U
Direct Labor Efficiency variance $316,800.00 F

Solution c:

Variable Overhead Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
52400 $149.43 $7,830,000.00 52400 $140.00 $7,336,000.00 59000 $140.00 $8,260,000.00
$494,000.00 U $924,000.00 F
Variable overhead rate variance Variable overhead efficiency variance
Variable overhead rate variance $494,000.00 U
Variable overhead efficiency variance $924,000.00 F